OpenStudy (anonymous):

Expand the series and evaluate:

OpenStudy (anonymous):

\[\sum_{k=1}^{4}(2k-1)\]

OpenStudy (anonymous):

a) 2 + 4 + 6 + 8 = 16 b) 1 + 3 + 5 + 7 = 16 c) 1 + 3 + 5 + 7 = 21 d) 2 + 4 + 6 + 8 = 20

hartnn (hartnn):

2k-1 plug in k= 1 to get 1st term 2(1) -1 = .... ? plug in k= 2 to get 2nd term 2(2) -1 = .... ? plug in k= 3 to get 3rd term 2(3) -1 = .... ? lastly plug in k= 4 to get 4th term 2(4) -1 = .... ?

OpenStudy (anonymous):

b) 1 + 3 + 5 + 7 = 16 thanks man !(:

hartnn (hartnn):

correct! welcome ^_^